# Characteristic Learning for Provable One Step Generation

### v3

@@ C.1. Auxiliary Properties
According to Lemma D.2, the following inequality holds
(F.24) $\inf_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)\leq\frac{CR^{2}}{N^{2}},$
where $C$ is a constant only depending on $d$ and $\sigma$. On the other hand, by applying Lemma A.12, the VC-dimension of this deep neural network class ${\mathscr{G}}$ is given as
(F.25) $\operatorname{VCdim}(\Pi_{k}{\mathscr{G}})\leq CN^{d+2}\log N,$
where $C$ is an absolute constant. Plugging (F.24) and (F.25) into Lemma F.2 yields
@@
Finally, we relate ${\mathcal{R}}_{T}(\widehat{g})$ to ${\mathcal{D}}(\widehat{g})$ by Lemma F.1. Finally, using Theorem 3.11 completes the proof. ∎
@@
(F.26) $u(t,s)=u(t_{k-1},t_{\ell-1})+\partial_{t}u(\theta_{k-1}^{t},t_{\ell-1})(t-t_{k-1})+\partial_{s}u(t_{k-1},\theta_{\ell-1}^{s})(s-t_{\ell-1}),$
@@
For $1\leq\ell=k\leq K$, integrating both sides of (F.26) on $(t,s)\in[t_{k-1},t_{k}]\times[t,t_{k}]$ yields
@@
(F.27) $\leq\sup_{(t,s)\in[0,T]^{2}}\big{\{}|\partial_{t}u(t,s)|+|\partial_{s}u(t,s)|\big{\}}\frac{T^{3}}{3K^{2}}.$
For $1\leq\ell<k\leq K$, integrating both sides of (F.26) on $(t,s)\in[t_{k-1},t_{k}]\times[t_{\ell-1},t_{\ell}]$ yields
@@
By a similar argument to (F.27), it follows that

### v5

@@ D.1. Auxiliary Properties
According to Lemma E.2, the following inequality holds
(G.24) $\inf_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)\leq\frac{CR^{2}}{N^{2}},$
where $C$ is a constant only depending on $d$ and $\sigma$. On the other hand, by applying Lemma B.12, the VC-dimension of this deep neural network class $\mathscr{G}$} is given as
(G.25) $\operatorname{VCdim}(\Pi_{k}{\mathscr{G}})\leq CN^{d+2}\log N,$
where $C$ is an absolute constant. Plugging (G.24) and (G.25) into Lemma G.2 yields
@@
[…]
@@
(G.26) $u(t,s)=u(t_{k-1},t_{\ell-1})+\partial_{t}u(\theta_{k-1}^{t},t_{\ell-1})(t-t_{k-1})+\partial_{s}u(t_{k-1},\theta_{\ell-1}^{s})(s-t_{\ell-1}),$
@@
For $1\leq\ell=k\leq K$, integrating both sides of (G.26) on $(t,s)\in[t_{k-1},t_{k}]\times[t,t_{k}]$ yields
@@
(G.27) $\leq\sup_{(t,s)\in[0,T]^{2}}\big\{|\partial_{t}u(t,s)|+|\partial_{s}u(t,s)|\big\}\frac{T^{3}}{3K^{2}}.$
For $1\leq\ell<k\leq K$, integrating both sides of (G.26) on $(t,s)\in[t_{k-1},t_{k}]\times[t_{\ell-1},t_{\ell}]$ yields
@@
By a similar argument to (G.27), it follows that